Logiweb diagnose of Frozen

Up Help

No errors found

The pyk compiler, version 0.grue.20060417+ by Klaus Grue GRD-2006-06-21.UTC:07:04:17.579248 = MJD-53907.TAI:07:04:50.579248 = LGT-4657590290579248e-6